OpenStudy (anonymous):

can someone help my find the limit if x->0 sinx/ cubed root of x

hartnn (hartnn):

multiply and divide by x , to get the function in the form of sin x/ x

OpenStudy (anonymous):

can u draw it out so i can see what ur talking about

hartnn (hartnn):

\(\huge\frac{sin x}{x}\:\frac{x}{\sqrt[3]{x}}\)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

and then from there which method would u use

hartnn (hartnn):

u know the formula : \[\lim_{x \rightarrow 0}\: sinx / x = 1\] ? use that for first fraction.

OpenStudy (anonymous):

oh ok! omg now it makes sense thank you @hartnn

hartnn (hartnn):

welcome :) what did u get the final answer as ?

OpenStudy (anonymous):

hang on lolz

OpenStudy (anonymous):

is it 0?

hartnn (hartnn):

yup, its 0 good work :)

OpenStudy (anonymous):

yay!!!

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

i need help again actually with number 84 now

hartnn (hartnn):

ok, if f(x) = root x what will be f(x+h) ?? u know?

OpenStudy (anonymous):

im confused on tht part

hartnn (hartnn):

ok, to get f(x+h) just replace x by (x+h) in f(x) so f(x+h) will be \(\sqrt{x+h}\) ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and then what?

hartnn (hartnn):

i m using h instead of delta x just for convenience here. so now u put that in {f(x+h)-f(x) }/ h to get \(\frac{\sqrt{x+h}-\sqrt{h}}{h}\) got this ?

OpenStudy (anonymous):

yeah i got tht part

hartnn (hartnn):

now mutiply and divide by \(\sqrt{x+h}+\sqrt{h}\) and use (a+b)(a-b) = a^2 - b^2 and tell me what u get in numerator ?

OpenStudy (anonymous):

oh so ur multiplying it by its conjugate right ?

hartnn (hartnn):

yup, so that i get only h in numerator and that i can cancel it with h in denominator ok?

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

i made a typo ..... it should be \(-\sqrt{x} \) and not \(-\sqrt{h} \) in the numerator

OpenStudy (anonymous):

wait a minute now im confused a bit

hartnn (hartnn):

because itsf(x+h) - f(x) which is \(\sqrt{x+h} \) \(-\sqrt{x} \)

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

so now now mutiply and divide by \(\sqrt{x+h}+\sqrt{x}\) and use (a+b)(a-b) = a^2 - b^2 and tell me what u get in numerator ?

OpenStudy (anonymous):

would it still be h in the numerator i think..

hartnn (hartnn):

yup, because (x+h)- x = x+h-x = h this h cancels out with h in denominator and u have only 1 in numerator, right ? and what u have in denominator now?

OpenStudy (anonymous):

do you have \[\sqrt{x+h} + \sqrt{x}\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

idk lolz

hartnn (hartnn):

yes, u are right, be confident :) now just put h =0 because lim h-> 0 and tell me what u get in denominator after putting h=0

OpenStudy (anonymous):

\[\sqrt{x+0} +\sqrt{x}\]

OpenStudy (anonymous):

is it?

hartnn (hartnn):

yup, thats your denominator = \(2\sqrt{x}\) and then your final answer would be \(\huge\frac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):

hang on im confused how do u get \[\frac{ 1 }{ 2\sqrt{x} }\]

hartnn (hartnn):

\(\sqrt{x+0} +\sqrt{x}=\sqrt{x} +\sqrt{x}=2\sqrt{x}\) and there was 1 in the numerator, previously, right ?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thanks!

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

u r a very helpful math teacher!